# FlexDiT: Dynamic Token Density Control for Diffusion Transformer

### v1

@@ 7 Captions in Figure 5
8 Additional visualization
We provide additional visualizations at a resolution of $512\times 512$ on FlexDiT-XL from Figure 6 to Figure 17. The class labels, including “arctic wolf”, “volcano”, “cliff drop-off”, “balloon”, “sulphur-crested cockatoo”, “lion”, “otter”, “coral reef”, “macaw”, “red panda”, ”“husky”, and “panda”, correspond to the same cases presented in DiT. Readers can compare our results with those of DiT. Our samples demonstrate comparable image quality and fidelity. The classifier-free guidance scale is set to 4.0, and all samples shown here are uncurated.
Figure 6: Uncurated $512\times 512$ FlexDiT-XL samples.
@@ 8 Additional visualization
Class label = “arctic wolf” (270)
Figure 7: Uncurated $512\times 512$ FlexDiT-XL samples.
@@
Class label = “volcano” (980)
Figure 8: Uncurated $512\times 512$ FlexDiT-XL samples.
@@
Class label = “husky” (250)
Figure 9: Uncurated $512\times 512$ FlexDiT-XL samples.
@@
Class label = “sulphur-crested cockatoo” (89)
Figure 10: Uncurated $512\times 512$ FlexDiT-XL samples.
@@
Class label = “cliff drop-off” (972)
Figure 11: Uncurated $512\times 512$ FlexDiT-XL samples.
@@
Class label = “balloon” (417)
Figure 12: Uncurated $512\times 512$ FlexDiT-XL samples.
@@
Class label = “lion” (291)
Figure 13: Uncurated $512\times 512$ FlexDiT-XL samples.
@@
Class label = “otter” (360)
Figure 14: Uncurated $512\times 512$ FlexDiT-XL samples.
@@
Class label = “red panda” (387)
Figure 15: Uncurated $512\times 512$ FlexDiT-XL samples.
@@
Class label = “panda” (388)
Figure 16: Uncurated $512\times 512$ FlexDiT-XL samples.
@@
Class label = “coral reef” (973)
Figure 17: Uncurated $512\times 512$ FlexDiT-XL samples.
@@
Class label = “macaw” (88)

### v2

@@ A.4 Captions in Figure 5
Model Method FID Speed-up
DiT-B ToMeSD 29.24 +20%
Ours 8.23 +68%
DiT-XL ToMeSD 14.74 +66%
Ours 2.38 +87%
Table 16: Comparison of our method with ToMeSD on DiT models.
Model Resolution Method FLOPs FID IS Speed-up
DiT-XL $512\times 512$ DyDiT 375 (-29%) 2.88(-0.16) - +31%
Ditfastattn - 4.52(+1.36) 180.34 +98%
Ours 286 (-46%) 2.96(-0.08) 242.4 +145%
DiT-XL $256\times 256$ TokenCache 72.25(-39%) 2.37 (+0.13) 262.00 +32%
Ours 68.05(-43%) 2.38 (+0.11) 276.39 +87%
Table 17: Comparison of our method with DyDiT on DiT models.
Method FID Flops Interations
DiT-XL 19.47 118.64 400K
Ours (from scratch) 15.11 68.05 400K
Table 18: Comparison of our method with ToMeSD on DiT models.
A.5 Additional visualization
We provide additional visualizations at a resolution of $512\times 512$ on SparseDiT-XL from Figure 6 to Figure 17. The class labels, including “arctic wolf", “volcano", “cliff drop-off", “balloon", “sulphur-crested cockatoo", “lion", “otter", “coral reef", “macaw", “red panda", "“husky", and “panda", correspond to the same cases presented in DiT. Readers can compare our results with those of DiT. Our samples demonstrate comparable image quality and fidelity. The classifier-free guidance scale is set to 4.0, and all samples shown here are uncurated.
Figure 6: Uncurated $512\times 512$ SparseDiT-XL samples.
@@ A.5 Additional visualization
Class label = “arctic wolf" (270)
Figure 7: Uncurated $512\times 512$ SparseDiT-XL samples.
@@
Class label = “volcano" (980)
Figure 8: Uncurated $512\times 512$ SparseDiT-XL samples.
@@
Class label = “husky" (250)
Figure 9: Uncurated $512\times 512$ SparseDiT-XL samples.
@@
Class label = “sulphur-crested cockatoo" (89)
Figure 10: Uncurated $512\times 512$ SparseDiT-XL samples.
@@
Class label = “cliff drop-off" (972)
Figure 11: Uncurated $512\times 512$ SparseDiT-XL samples.
@@
Class label = “balloon" (417)
Figure 12: Uncurated $512\times 512$ SparseDiT-XL samples.
@@
Class label = “lion" (291)
Figure 13: Uncurated $512\times 512$ SparseDiT-XL samples.
@@
Class label = “otter" (360)
Figure 14: Uncurated $512\times 512$ SparseDiT-XL samples.
@@
Class label = “red panda" (387)
Figure 15: Uncurated $512\times 512$ SparseDiT-XL samples.
@@
Class label = “panda" (388)
Figure 16: Uncurated $512\times 512$ SparseDiT-XL samples.
@@
Class label = “coral reef" (973)
Figure 17: Uncurated $512\times 512$ SparseDiT-XL samples.
@@
Class label = “macaw" (88)